Please help me!!! I am not sure if I have set this up properly!
a dehydrated patient needs a 3.24% saline IV. Unfortunately the hospital only has bags of 1% and 9% saline solutions. How many liters of each of these solutions should be mixed together to yield 5 liters of the desired concentration. This is what I have done please let me know if I've done it right or give me guidance to the right way.

x + y = 5; x = 5 - y
x + 9y = 5 x 3.24 = 16.2
5 - y + 9y = 16.2
8y = 11.2
y = 1.4 L of 9%
x = 3.6 L of 1%

Answer by stanbon(75887)

dehydrated patient needs a 3.24% saline IV. Unfortunately the hospital only has bags of 1% and 9% saline solutions. How many liters of each of these solutions should be mixed together to yield 5 liters of the desired concentration.
------
Equation:
salt + salt = salt
0.01x + 0.09(5-x) = 0.0324*5
--------
Multiply thru by 10000 to get:
100x + 900*5 - 900x = 324*5
-----
-800x = -576*5
x = 3.6 liters (amt. of 1% solution needed)
5-x = 1.4 liters (amt. of 9% solution needed)
==================================================
Congratulations; you have the correct answers.
==================================================
Cheers,
Stan H.
============